A father is 5 times as old as his son. In 18 years, the father will be twice
as old as his son. How old is the father at present?

Answer by rothauserc(4718)

let x be the age of the father's son
:
father's age = 5x
:
5x+18 = 2 * (x+18)
:
5x+18 = 2x+36
:
3x = 18
:
x = 6
:
*******************************************
Father's current age = 5 * 6 = 30 years old
*******************************************
:
check the answer
:
30+18 = 2 * (6+18)
:
48 = 48
:
answer checks